# NativQA Framework: Enabling LLMs with Native, Local, and Everyday Knowledge

### v1

@@ Abstract
Firoj Alam,1 Md Arid Hasan,2 Sahinur Rahman Laskar,3 Mucahid Kutlu,4 Shammur Absar Chowdhury1 1Qatar Computing Research Institute, Qatar, 2University of New Brunswick, Canada, 3UPES, India, 4Qatar University, Qatar fialam@hbku.edu.qa, arid.hasan@unb.ca https://gitlab.com/nativqa/nativqa-framework
@@
Firoj Alam,1 Md Arid Hasan,2 Sahinur Rahman Laskar,3 Mucahid Kutlu,4 Shammur Absar Chowdhury1 1Qatar Computing Research Institute, Qatar, 2University of New Brunswick, Canada, 3UPES, India, 4Qatar University, Qatar fialam@hbku.edu.qa, arid.hasan@unb.ca https://gitlab.com/nativqa/nativqa-framework
@@ 1 Introduction
Figure 1: NativQA  framework has already been used to collect QA pairs from 39 locations, 24 countries and covering 7 languages. (a) The plot shows city-level data distribution grouped by language. The inner circle represents languages, and the outer wedges represent cities (with country names). The size of each wedge corresponds to the number of samples. (b) topics that has been evaluated with the associated datasets.
@@
Large Language Models (LLMs) have exhibited cultural bias toward certain cultures, largely due to the predominance of training data in high-resource languages and the limited digital representation of low-resource languages Li et al. (2024a); Pawar et al. (2024); Durmus et al. (2024). Given that most LLMs are trained primarily on English or translated data, which often fails to capture how people naturally speak, ask questions, and express their needs in their native languages and dialects. Since the emergence of LLMs, there has been growing concern about how to develop and benchmark models that perform equitably across diverse languages and cultures Li et al. (2024b); Myung et al. (2024); Kim et al. (2024); Pawar et al. (2024). In addition to cultural representation, [^1] Cultural ability includes understanding the diversity of cultural elements across different cultures Pawar et al. (2024). In the context of cultural QA, it can be defined as question answering that reflects the types of questions likely to be asked and answered by people of a particular culture, either in their native or second language Arora et al. (2024). there is also concern about whether LLMs can effectively respond to everyday, long-form, and complex queries that require paragraph-length answers Arora et al. (2024); Myung et al. (2024); Hasan et al. (2024). These concerns have driven efforts to develop multilingual Üstün et al. (2024); Touvron et al. (2023); Team et al. (2024) and language-specific Team et al. (2025); Nahin et al. (2025); Nguyen et al. (2024) closed and open-source LLMs. To benchmark and fine-tune models with region- and culture-specific knowledge, there have been efforts to develop language- and culture-specific resources, along with benchmarked and fine-tuned models Chiu et al. (2024); Shi et al. (2024); Li et al. (2024a). Methodologies involved in these efforts include collecting QA pairs from community question-answering forums Shi et al. (2024), hiring native speakers to write questions and answers Chiu et al. (2024), translating content from English into local languages Myung et al. (2024), and using LLMs to generate QA pairs Putri et al. (2024). A common finding across studies is that while most LLMs perform reasonably well on high-resource languages, they tend to struggle with mid- and low-resource languages Myung et al. (2024). Several studies have demonstrated that fine-tuning with culture-specific knowledge can improve model performance, even in smaller or quantized models Li et al. (2024a, b); Hasan et al. (2024). These findings suggest that achieving inclusivity in LLMs requires datasets that reflect diverse regions, cultures, and content generated by native users. One of the key challenges in creating linguistically and culturally rich datasets is the substantial manual effort they require. As LLMs evolve, there is a growing need for scalable tools to build culturally and geographically relevant datasets. However, current resources often lack support for location- and language-specific development, limiting broader inclusivity and applicability. To bridge this gap, we propose NativQA framework–a system that enables communities to easily build culture-specific QA datasets for enriching and benchmarking LLMs. The NativQA framework takes a seed query (e.g., “visit Baladna Farm in Qatar”) and a location specifier (e.g., “qa”) as inputs, and generates a list of QA pairs as output, as illustrated in Figure 2. We evaluated NativQA framework for 39 locations, from 24 countries covering 7 languages and 18 topics, which resulted over 300K data points as shown in Figure A2. To summarize, the contributions of our study are as follows.
@@
  • We propose the NativQA framework for creating culture- and region-specific natural question-answering datasets. This framework aims to promote inclusivity in LLMs by empowering the community to develop diverse, culturally aligned QA resources and comprehensive benchmark datasets.
@@
  • NativQA framework offers several key features: (i) seamless integration of both user- and LLM-generated queries to collect QA pairs; (ii) location-agnostic – users can collect data from any region; (iii) support for multiple search engines; (iv) image search capability; (v) automated removal of duplicate QA pairs; (vi) domain reliability checks; and (vii) an efficient caching mechanism to reduce redundant API calls, leading to cost savings and mitigating time-out issues.
@@
  • We conducted an extensive evaluation of NativQA framework to assess its feasibility for broader adoption and demonstrated that it can be effectively used across different locations, cultures, and languages.
@@ 2 NativQA framework
In Figure 2, we present the NativQA framework consisting of three inter-connected modules such as (i) Query Collection, (ii) QA Collection and (iii) QA Validation. The top part of the figure shows the entire process while the bottom part shows an example from query collection to QA validation. In the below sections we discuss them in details.
@@ 2.1 Query Collection (QC) Module
The objective of this module is to collect open-ended queries focused on various predetermined topics derived from common concepts in everyday communication. The set of topics should be established prior to beginning the query collection process, as this facilitates the gathering of topic-specific queries. As shown in Figure A2, the framework has been evaluated using a predetermined set of 18 topics that are culture- or region-dependent (e.g., Events, Literature, etc.). However, the topics can be anything and can be task specific. The query collection process can be carried out using three different approaches:
@@
The manual query collection process requires recruiting location- or region-specific annotators. Moreover, like any other annotation task, it is time-consuming and requires monitoring for both progress and quality. For example, in Hasan et al. (2024), country-specific annotators were hired to manually write topic specific queries. In contrast, the latter approaches are more cost- and time-effective. The template-based approach involves defining location-agnostic templates, which can be seamlessly adapted for different locations. The LLM-based approach requires prompt engineering to generate appropriate queries. We use the term seed query ($Q_{0}$) to refer to the queries generated by all approaches. These queries are then used in later steps by the QA collection module to gather QA pairs and additional queries. Filtering: Given that queries can have exact and/or near duplicates, it is important to remove such duplicates as part of the filtering process. The filtering module handles this by using exact matching for exact duplicates and a similarity-based approach for near duplicates. These steps results in the final set of queries.
@@ 2.2 QA Collection (QAC) Module
The next step is to collect QA pairs that potentially cover the topics represented by a search engine, e.g., Google. Most search engines support features like “People also ask” or “Related queries,” which list questions asked by real users that are potentially relevant to the initial seed query, as shown in Figure A1. Moreover, these questions are accompanied by answers extracted by the search engine, along with links to the sources of the answers. The QA collection module implements Algorithm 1 (outlined in Appendix). It takes the seed queries $Q_{0}$ and the number of iterations $N_{\text{iter}}$ as input. For each iteration $i\in N_{\text{iter}}$, it collects QA pairs $P_{QA}^{i}$ and related queries $S^{i}_{\text{rel}}$ for each query $q\in Q$, then passes them to the filtering module and updates the current query set $Q$. This process is repeated for all iterations to obtain the final QA set, $S_{QA}$, for the enriched queries $Q$. The idea of the filtering module is to remove duplicates that may arise in different iterations. The QA collection module leverages search engine APIs to collect QA pairs, allowing location and language to be specified in the API call. These parameters define the search criteria. Location parameter is especially important, as the focus is on collecting QA pairs from specific regions. The module also stores relevant metadata (e.g., answer source URL) for later use.
@@ 2.3 QA Validation (QAV) Module
The next step of the framework is to validate the QA pairs. It is important for several reasons: (i) the retrieved related queries might be less relevant, (ii) answers might be incomplete, (iii) the source of the answer can be less reliable. Therefore, we opted to validate the QA pairs using the following two approaches.
Domain Reliability Checking (DRC).
The answers collected by NativQA framework include a link to the source web page from which the answer was extracted. Therefore, the aim of the DRC approach is to retain QA pairs based on the reliability of the web domain where the answer appears. We hypothesize that answers from pages hosted on reliable domains are more likely to be trustworthy and can be directly included in the dataset. NativQA framework currently supports two options for DRC:
@@
  • In the current implementation, NativQA framework provides $2,080$ source URLs obtained from the MultiNativQA dataset Hasan et al. (2024), where annotators identified them as very reliable.
This approach is more practical and scalable for developing large-scale fine- and instruction-tuning data, as it reduces the manual effort required to obtain hand-curated answers. However, due to the semi-supervised nature of the approach, it has certain limitations. Some domains are authentic (e.g., Google, Facebook); however, the content they host might not be reliable due to user-generated content of questionable quality.
@@
To increase more reliability of the curated questions with answers, different steps can be taken. Currently NativQA framework supports two different approaches:
  • Manual annotation: This approach involves manually reviewing and editing the answers for the QA pairs. It includes three types of annotations. (i) Question validation: Human annotators assess the quality of the questions by assigning one of the following labels: good or bad. A “good question” is defined as a fact-seeking question that can be answered with an entity or an explanation. In contrast, a “bad question” is either ambiguous, incomprehensible, based on a clear false presupposition, opinion-seeking, or does not seek factual information. (ii) Answer editing: If a given answer does not fully address the question, or includes extra or incorrect information, the annotator must edit the answer using content from the source web page to ensure completeness and accuracy. Annotators can only use the provided source web pages to maintain the scope and the reliability of the answers. (iii) Location relevance: This task involves determining whether the question is relevant to the specified [LOCATION].
  • LLM-based annotation: While manual annotation by humans is more reliable, it is also time-consuming and costly. To reduce such effort, NativQA framework supports the use of LLM-based approaches as models like GPT-4 and other large LLMs have demonstrated effectiveness in improving task-specific model performance Li et al. (2023, 2024c). In the LLM-based approach, a concise version of the annotation guidelines and task definitions is adapted into the prompt to check the question, edit the answer, and identify location relevance.
@@ 3 Evaluation of NativQA framework
Evaluated in multiple studies. The NativQA framework used in Hasan et al. (2024) to collect data from 9 locations in 7 languages, where seed queries and QA validation performed manually. This effort resulted in approximately 64K annotated QA pairs. Mousi et al. (2025) used the framework for Arabic cultural and dialectal QA data collection to benchmark the cultural and dialectal capabilities of LLMs. Evaluation through additional data collection. We further evaluated the framework for additional 30 locations by creating seed query templates. We developed Arabic and English seed query templates from the seed queries that were used to collect the MultiNativQA dataset Hasan et al. (2024). We selected seed queries that contain mentions of locations in the MultiNativQA, which were then used with the [LOCATION] template (e.g., women’s wedding attire in [LOCATION]). The [LOCATION] template was subsequently replaced with other locations to create seed queries for different regions. We manually validated the seed queries generated from the template and removed those that were specific to a particular location (e.g., 1971 liberation war in Bangladesh/[LOCATION]). For Arabic, we used only the seed queries from the template that were employed to collect QA data for Qatar, while the English queries were adopted from the manual seed queries for both Bangladesh and New York. In Figure A2, we report the distribution of the locations and languages for which data has been collected using the NativQA framework. The reported distribution includes datasets from MultiNativQA as well as additional datasets we collected to evaluate the efficacy of the framework. Capitalizing LLMs. In the NativQA framework, we used LLMs in two stages: (i) query generation and (ii) QA validation. Query generation. For query generation, we designed prompts specifying the location and topic, and used LLMs to generate queries. To assess the quality of the generated queries, we manually reviewed a random sample and found that they were reasonable and reflected the specified location and topic. Notably, most of them matched human-generated queries at the surface level. Generating a reasonable set of queries requires some prompt engineering. We will release the scripts with prompts used as part of the framework. QA validation. For QA validation, we manually evaluated the output of LLMs on 500 QA pairs collected from Egypt. We only selected due to the availability of annotators from Egypt. The goal was to assess whether human annotators preferred the LLM-edited answers or the ones retrieved from a search engine. We provided clear instructions to the annotators to select the better answer based on accuracy, clarity and readability, as well as completeness and relevance. Annotators were given three options: (i) Answer 1 (our original answer), (ii) LLM-edited answer, and (iii) neither. We recruited two annotators who are native Arabic speakers residing in Egypt. They were compensated at a standard rate managed by a third-party company. We computed inter-annotator agreement using observed agreement, Gwet’s AC1 Gwet (2008), and Cohen’s Kappa. The observed agreement was 0.842, indicating strong overall alignment. Gwet’s AC1 was 0.806, while Cohen’s Kappa was relatively low ($\kappa$ = 0.147), reflecting the known limitation of Kappa in skewed distributions—since annotators selected the LLM-edited answer in the majority of cases, Cohen’s method underestimates agreement when one class dominates. The high observed agreement suggests that using LLMs to edit answers may be a promising direction. For both query generation and QA validation we used GPT-4o (version: 2024-11-20). However, any models can be used for such tasks.

### v2

@@ Abstract
Firoj Alam,1 Md Arid Hasan,2 Sahinur Rahman Laskar,3 Mucahid Kutlu,4 Kareem Darwish,1 Shammur Absar Chowdhury1 1Qatar Computing Research Institute, Qatar, 2University of New Brunswick, Canada, 3UPES, India, 4Qatar University, Qatar fialam@hbku.edu.qa, arid.hasan@unb.ca https://gitlab.com/nativqa/nativqa-framework
@@
Firoj Alam,1 Md Arid Hasan,2 Sahinur Rahman Laskar,3 Mucahid Kutlu,4 Kareem Darwish,1 Shammur Absar Chowdhury1 1Qatar Computing Research Institute, Qatar, 2University of New Brunswick, Canada, 3UPES, India, 4Qatar University, Qatar fialam@hbku.edu.qa, arid.hasan@unb.ca https://gitlab.com/nativqa/nativqa-framework
@@ 1 Introduction
Figure 1: The NativQA framework has been used to collect QA pairs across 39 locations in 24 countries, covering 7 languages. (a) The plot shows city-level data distribution grouped by language—inner circles denote languages, outer wedges represent cities (with country names), sized by sample count. (b) Topics evaluated with the corresponding datasets are also shown.
@@
Large Language Models (LLMs) have exhibited cultural bias toward certain cultures, largely due to the predominance of training data in high-resource languages and the limited digital representation of low-resource languages Li et al. (2024a); Pawar et al. (2024); Durmus et al. (2024). Given that most LLMs are trained primarily on English or translated data, which often fails to capture how people naturally speak, ask questions, and express their needs in their native languages and dialects. Since the emergence of LLMs, there has been growing concern about how to develop and benchmark models that perform equitably across diverse languages and cultures Li et al. (2024b); Myung et al. (2024); Kim et al. (2024); Pawar et al. (2024). In addition to cultural representation, [^1] Cultural ability includes understanding the diversity of cultural elements across different cultures Pawar et al. (2024). In the context of cultural QA, it can be defined as question answering that reflects the types of questions likely to be asked and answered by people of a particular culture, either in their native or second language Arora et al. (2024). there is also concern about whether LLMs can effectively respond to everyday, long-form, and complex queries that require paragraph-length answers Arora et al. (2024); Myung et al. (2024); Hasan et al. (2024). These concerns have driven efforts to develop multilingual Üstün et al. (2024); Touvron et al. (2023); Team et al. (2024) and language-specific Team et al. (2025); Nahin et al. (2025); Nguyen et al. (2024) closed and open-source LLMs. To benchmark and fine-tune models with region- and culture-specific knowledge, there have been efforts to develop language- and culture-specific resources, along with benchmarked and fine-tuned models Chiu et al. (2024); Shi et al. (2024); Li et al. (2024a). Methodologies involved in these efforts include collecting QA pairs from community question-answering forums Shi et al. (2024), hiring native speakers to write questions and answers Chiu et al. (2024), translating content from English into local languages Myung et al. (2024), and using LLMs to generate QA pairs Putri et al. (2024). A common finding across studies is that while most LLMs perform reasonably well on high-resource languages, they tend to struggle with mid- and low-resource languages Myung et al. (2024). Several studies have demonstrated that fine-tuning with culture-specific knowledge can improve model performance, even in smaller or quantized models Li et al. (2024a, b); Hasan et al. (2024). These findings suggest that achieving inclusivity in LLMs requires datasets that reflect diverse regions, cultures, and content generated by native users. One of the major challenges in creating linguistically and culturally rich datasets is the significant manual effort involved. As LLMs continue to advance, the need for scalable methods to build these resources becomes more critical. However, there is currently a lack of tools designed to support the efficient development of such datasets. To bridge this gap, we propose NativQA framework–a system that enables communities to easily build culture-specific QA datasets for enriching and benchmarking LLMs. We evaluated NativQA framework for 39 locations, from 24 countries covering 7 languages and 18 topics, which resulted over 300K data points as shown in Figure 1. To summarize, the contributions of our study are as follows.
@@
  • We propose NativQA framework, for creating culture- and region-specific QA datasets that promote inclusivity in LLMs by enabling community-driven development of diverse, culturally aligned benchmarks.
@@
  • The NativQA framework offers several key features: (i) seamless integration of both user- and LLM-generated queries to collect QA pairs; (ii) location-agnostic – users can collect data from any region; (iii) support for multiple search engines; (iv) image search capability; (v) automated removal of duplicate QA pairs; (vi) domain reliability checks; and (vii) an efficient caching mechanism to reduce redundant API calls, leading to cost savings and mitigating time-out issues.
@@
  • We conducted an extensive evaluation to assess its feasibility for broader adoption and demonstrated that it can be effectively used across different locations, cultures, and languages.
@@ 2 NativQA framework
In Figure 2, we present the NativQA framework consisting of three inter-connected modules such as (i) Query Collection, (ii) QA Collection, and (iii) QA Validation. The top part of the figure shows the entire process while the bottom part shows an example from query collection to QA validation. In the below sections we discuss them in details.
@@ 2.1 Query Collection (QC) Module
The objective of this module is to collect open-ended queries focused on various predetermined topics derived from common knowledge in everyday communication. The set of topics should be established prior to beginning the query collection process, as this facilitates the gathering of topic-specific queries. As shown in Figure 1, the framework has been evaluated using a predetermined set of 18 topics that are culture- or region-dependent (e.g., Events, Literature, etc.). However, the topics can be anything and can be task specific. The query collection process can be carried out using three different approaches:
@@
The manual query collection process requires recruiting location- or region-specific annotators. Moreover, like any other annotation task, it is time-consuming. For example, in Hasan et al. (2024), country-specific annotators were hired to manually write topic specific queries. In contrast, the latter approaches are more cost- and time-effective. The template-based approach involves defining location-agnostic templates, which can be seamlessly adapted for different locations. The LLM-based approach requires prompt engineering to generate appropriate queries. We use the term seed query ($Q_{0}$) to refer to the queries generated by all approaches. These queries are then used in later steps by the QA collection module to gather QA pairs and additional queries. Filtering: Given that queries can have exact and/or near duplicates, it is important to remove such duplicates as part of the filtering process. The filtering module handles this by using exact matching for exact duplicates and a similarity-based approach for near duplicates. These steps results in the final set of queries.
@@ 2.2 QA Collection (QAC) Module
The next step is to collect QA pairs using a search engine, e.g., Google. Most search engines support features like “People also ask” or “Related queries,” which list questions asked by real users that are potentially relevant to the initial seed query, as shown in Figure A1. Moreover, these questions are accompanied by answers extracted by the search engine, along with links to the sources of the answers. The QA collection module implements Algorithm 1 (outlined in Appendix). It takes the seed queries $Q_{0}$ and the number of iterations $N_{\text{iter}}$ as input. For each iteration $i\in N_{\text{iter}}$, it collects QA pairs $P_{QA}^{i}$ and related queries $S^{i}_{\text{rel}}$ for each query $q\in Q$, then passes them to the filtering module and updates the current query set $Q$. This process is repeated for all iterations to obtain the final QA set, $S_{QA}$, for the enriched queries $Q$. The idea of the filtering module is to remove duplicates that may arise in different iterations. The QA collection module leverages search engine APIs to collect QA pairs, specifying location and language parameters in the API call. These parameters define the search criteria. Location parameter is especially important, as the focus is on collecting QA pairs from specific regions. The module also stores relevant metadata (e.g., answer source URL) for later use.
@@ 2.3 QA Validation (QAV) Module
The next step of the framework is to validate the QA pairs. It is important for several reasons: (i) the retrieved related queries might be less relevant, (ii) answers might be incomplete, (iii) the source of the answer can be less reliable. For the QAV the framework integrates the below two approaches.
1. Domain Reliability Checking (DRC).
The answers collected by NativQA framework include a link to the source web page from which the answer was extracted. Therefore, the aim of the DRC is to retain QA pairs based on the reliable web domain where the answer appears. We hypothesize that answers from pages hosted on reliable domains are more likely to be trustworthy and can be directly included in the dataset. The framework currently supports two options for DRC:
@@
  • In the current implementation, NativQA framework provides $2,080$ source URLs obtained from the MultiNativQA dataset Hasan et al. (2024), where annotators identified them as very reliable. These domain URLs used to verify the reliability of the answers.
This approach is more practical and scalable for developing large-scale QA pairs, as it reduces the manual effort required to obtain hand-curated answers. However, due to the semi-supervised nature of the approach, it has certain limitations. Some domains are authentic (e.g., Google, Facebook); however, the content they host might not be reliable due to the questionable user-generated content.
@@
To increase more reliability of the curated QA pairs, different steps can be taken. Currently NativQA framework supports two different approaches:
  • Manual annotation: This approach involves manually reviewing and editing the answers. It includes three types of annotations. (i) Question validation: Human annotators assess the quality of the questions by assigning one of the following labels: good or bad. A “good question” is defined as a fact-seeking question that can be answered with an entity or an explanation. In contrast, a “bad question” is either ambiguous, incomprehensible, based on a clear false presupposition, opinion-seeking, or does not seek factual information. (ii) Answer editing: If a given answer does not fully address the question, or includes extra or incorrect information, the annotator must edit the answer using content from the source web page to ensure completeness and accuracy. Annotators can only use the provided source web pages to maintain the scope and the reliability of the answers. (iii) Location relevance: This task involves determining whether the question is relevant to the specified [LOCATION].
  • LLM-based annotation: While manual annotation by humans is more reliable, it is also time-consuming and costly. To reduce such effort, NativQA framework supports the use of LLM-based approaches as models like GPT-4 have demonstrated effectiveness in improving task-specific model performance Li et al. (2023, 2024c). In the LLM-based approach, a concise version of the annotation guidelines Hasan et al. (2024) and task definitions is adapted into the prompt to check the question, edit the answer, and identify location relevance.
@@ 3 Evaluation of NativQA framework
Evaluated in multiple studies. The NativQA framework used in Hasan et al. (2024) to collect data from 9 locations in 7 languages, where seed queries and QA validation performed manually. This effort resulted in approximately 64K annotated QA pairs. Mousi et al. (2025) used the framework for Arabic cultural and dialectal QA data collection to benchmark the cultural and dialectal capabilities of LLMs. Evaluation through additional data collection. We further evaluated the framework for additional 30 locations by creating seed query templates. We developed Arabic and English seed query templates from the seed queries that were used to collect the MultiNativQA dataset Hasan et al. (2024). We selected seed queries that contain mentions of locations in the MultiNativQA, which were then used with the [LOCATION] template (e.g., women’s wedding attire in [LOCATION]). The [LOCATION] template was subsequently replaced with other locations to create seed queries for different regions. We manually validated the seed queries generated from the template and removed those that were specific to a particular location (e.g., 1971 liberation war in Bangladesh/[LOCATION]). For Arabic, we used only the seed queries from the template that were employed to collect QA data for Qatar, while the English queries were adopted from the manual seed queries for both Bangladesh and New York. In Appendix C, we discuss the details of the data distribution (over 300K) that have been collected using the NativQA framework. Capitalizing LLMs. In the NativQA framework, we used LLMs in two stages: (i) query generation and (ii) LLM-based QA annotation. The use of LLMs for such tasks involves prompting them to produce the desired outcomes. We will release the prompts and scripts as part of the framework. LLM-based Query generation. For query generation, we designed prompts specifying the location and topic, and used LLMs to generate queries. To assess the quality of the generated queries, we manually reviewed a random sample and found that they were acceptable to human annotators (as they matched human-generated queries) and reflected the specified location and topic. LLM-based QA annotation. For LLM-based QA annotation, a concise version of the annotation guidelines and task definitions is adapted to the prompt to validate the question, edit the answer, and identify the location relevance. To manually assess the quality of the annotated QA pairs, we choose two different approaches: (i) QA validation by selecting the best answer, and (ii) QA annotation for LLM-based QA annotation.
  • Manual QA validation - selecting best answer. For QA validation for LLM-based QA annotation, we manually evaluated the output of LLMs on 500 QA pairs collected from Egypt. We selected this set due to the availability of annotators from Egypt. The goal was to assess whether human annotators preferred the LLM-edited answers or the ones retrieved from a search engine. We provided clear instructions to the annotators to select the better answer based on accuracy, clarity and readability, as well as completeness and relevance (see in Appendix D.1). Annotators were given three options: (i) Answer 1 (original answer), (ii) LLM-edited answer, and (iii) neither. We recruited two annotators who are native Arabic speakers residing in Egypt. They were compensated at a standard rate managed by a third-party company. We computed inter-annotator agreement using observed agreement and Gwet’s AC1 Gwet (2008), and Cohen’s Kappa. The observed agreement and Gwet’s AC1 were 0.842 and 0.806, respectively, indicating strong overall alignment. The high observed agreement suggests that using LLMs to edit answers may be a promising direction.
  • Manual QA validation - answer scoring. This approach involves manually checking the quality of LLM-edited answers and scoring them. We used a 5-point Likert scale for various evaluation metrics, including clarity, faithfulness, informativeness, and plausibility. These metrics are selected from various related studies on the evaluation of natural language explanations Huang et al. (2024); Zavolokina et al. (2024). We provide detailed annotation guidelines for human evaluators in Appendix D.2. Due to the availability of native human annotators, the evaluation was carried out for two locations, such as Sudan and Yemen, on $\sim$350 examples from each test set. Each LLM-edited answer was annotated by three and two human evaluators for Sudan and Yemen, respectively. We presented the average Likert-scale scores for all the evaluation metrics in Table 1. Furthermore, we measured annotation agreement on ordinal scales using the agreement index $r^{*}_{wg(j)}$ James et al. (1984), which evaluates how the observed rating variance compares to the maximum variance expected under total disagreement. The values over 0.85 for Sudan and 0.86 for Yemen indicate a strong agreement O’Neill (2017), as shown in Table 1. In addition, we analyze the annotations and find that 100% answers were not edited by annotators for the Sudan data. For Yemen, the percentage of not edited answer was 92%. Both human annotation results and analysis strongly suggest that LLMs are highly capable of QA annotation in the NativQA framework.
Location Clarity Faithfulness Informativeness Plausibility
Average Likert score
Khartoum, Sudan 4.41 4.42 4.26 4.29
Sanaa, Yemen 4.05 4.05 3.98 4.06
Agreement index $r^{*}_{wg(j)}$
Khartoum, Sudan 0.85 0.85 0.85 0.85
Sanaa, Yemen 0.88 0.86 0.89 0.88
Table 1: Average Likert scale and agreement scores for each human evaluation metric for LLM-edited QA.
For both query generation and LLM-based QA annotation, we used GPT-4o (version: 2024-11-20). However, any models can be used for such tasks.